OpenStudy (anonymous):

When \frac{8668}{25}+\frac{4141}{9}-\frac{5533}{25} is computed and written as a mixed number in simplest form, what is the fractional part of that mixed number?

OpenStudy (anonymous):

\[\frac{8668}{25}+\frac{4141}{9}-\frac{5533}{25}=\frac{26348}{45}=585+\frac{23}{45} \]